# Title keyword 5 for citation Title for citation Title for citation

### v2

@@ 8.4 Requirement of Real-Time AI
The 6G aims to achieve uRLLC to support various real-time applications, including autonomous driving, telemedicine, and industrial automation. These scenarios impose unprecedented requirements on network real-time responsiveness and reliability, necessitating data transmission and processing to be completed within milliseconds or even microseconds. However, the current state-of-the-art AI models, characterized by large parameter spaces and intensive computation, typically employ distributed deployment, which inevitably introduces additional latency. The computational complexity of AI models presents a direct challenge to the real-time capabilities of 6G networks. The inference process of deep learning models involves extensive matrix operations, especially when processing high-dimensional data, significantly increasing computational load. Achieving rapid model inference without sacrificing performance is a critical barrier to realizing the uRLLC objectives of 6G. The importance of real-time processing capabilities in 6G networks is not only reflected in enhancing user experience. Ensuring the regular operation and safety of critical services is crucial. For instance, the latency requirements for perception and decision-making systems in autonomous driving are extremely stringent; any delay could result in untimely system responses, thus posing safety risks. Therefore, real-time performance is a fundamental metric within 6G networks.
The challenge of model optimization is presented to achieve real-time AI in a 6G environment. Reducing the computational requirements of models without significantly compromising accuracy is possible through techniques such as model pruning, parameter quantization, and adopting more efficient network architectures, such as lightweight versions of CNN and RNN [308]. Additionally, adopting edge computing can reduce latency by shifting data processing tasks from the central cloud to the network edge [309]. Innovations in network architecture, such as network slicing technology, can also provide customized network resources for different services and applications, ensuring critical tasks receive the necessary bandwidth and latency guarantees [310]. Moreover, AI can manage network resources dynamically and accommodate real-time changes in network conditions. The development of hardware acceleration is also crucial for achieving real-time AI. Dedicated AI processors, such as Graphics Processing Units (GPU), Tensor Processing Units, and Field-Programmable Gate Arrays, can expedite the inference process of models, helping to meet the stringent real-time requirements of 6G networks.
@@ 8.5 Security and Privacy
The integration of AI and 6G introduces complex security and privacy challenges. The convergence of AI and 6G introduces security vulnerabilities at multiple levels. AI algorithms are susceptible to adversarial attacks, where slight, often imperceptible, perturbations in the input data can lead to incorrect outputs, thereby compromising the system’s integrity. Moreover, the reliance on large-scale data for training AI models poses privacy risks, as sensitive information may be inferred from the training data. Additionally, the distributed nature of 6G networks, with many devices connected at the edge, expands the attack surface, making it more challenging to secure against unauthorized access and data breaches [311].
Some existing work has inspired addressing these security and privacy issues. One aspect involves the development of robust AI models through techniques like adversarial training, which can enhance the resilience of AI systems against malicious to mitigate these inputs [312]. Another approach is the implementation of FL, where AI models are trained across multiple decentralized devices without exchanging the raw data, thereby preserving privacy [313]. In addition to technological solutions, regulatory frameworks and policies must be established to define clear data protection and cybersecurity standards in the context of AI-6G integration. This includes setting guidelines for responsible AI usage, ensuring compliance with privacy regulations such as the General Data Protection Regulation, and fostering international collaboration to combat cyber threats. Moreover, secure multi-party computation and homomorphic encryption can perform computations on encrypted data, ensuring data privacy even when processed by third parties [314].
@@ 8.6 Human-AI Interactive Collaboration
Integrating AI with 6G communication networks presents transformative opportunities for enhancing human-AI interactive collaboration. However, this convergence also introduces significant challenges that must be addressed to realize its full potential. One of the primary issues is the stringent real-time response requirements imposed by 6G’s ultra-low latency capabilities. Traditional AI algorithms, which often require substantial computational resources and time, may struggle to develop this, so developing lightweight AI models optimized for edge computing is essential. These models must be capable of rapid processing and decision-making, reducing the dependency on centralized data centers and minimizing latency. Furthermore, 6G networks will facilitate the transmission of diverse data types, including video, audio, text, and sensor data, necessitating human-AI interactive collaboration systems to handle and interpret multimodal data adeptly. Advances in deep learning and multimodal learning technologies are crucial for effectively fusing and understanding these varied data streams, thereby enhancing the naturalness and intuitiveness of interactions.
@@ 9 Future Works
The advent of 6G networks heralds a paradigm shift in managing network operations, with AI playing a central role. Adaptive learning mechanisms, leveraging AI, are set to become the cornerstone of smart network management. These mechanisms must handle the complexity and dynamism of 6G environments characterized by heterogeneous devices, variable traffic patterns, and the need for real-time processing.
Research is expected to focus on developing machine learning models that can continuously learn and adapt without human intervention. These models should be scalable to manage the exponential increase in connected devices and robust to ensure consistent performance under varying network states. One promising approach is using DRL, which can dynamically optimize network decisions. DRL can adjust network parameters in real-time, ensuring optimal performance for changing network loads and user mobility patterns. Moreover, integrating transfer learning techniques can enable these AI models to apply knowledge gained in one domain to new but related problems. This is particularly relevant in 6G networks that will operate in diverse environments and must quickly adapt to new conditions. Transfer learning can reduce the need for extensive retraining of models, thereby saving computational resources and time. Another focus area can be the development of self-healing networks using AI. By incorporating predictive analytics, networks can anticipate failures and automatically reroute traffic or adjust parameters to mitigate the impact. This proactive approach to fault management will be critical for maintaining the high reliability and availability required by 6G applications.
@@
The underlying network must be optimized to cater to their requirements for AI applications to reach their full potential. 6G networks, with their promise of higher bandwidths and lower latencies, provide an ideal platform for AI applications but also pose significant challenges in terms of architectural design.
Future research can delve into creating network architectures that are inherently flexible and can dynamically adapt to the needs of AI applications. Such architectures must support high data rates and the ability to process large volumes of data with minimal latency. AI can significantly enhance the capability of the network to manage these requirements by predicting traffic patterns and adjusting resources accordingly. A key component of such architectures is edge computing, which brings computational resources closer to the data sources. This is crucial for latency-sensitive AI applications that require near-instantaneous decision-making. Research should explore the development of edge computing solutions that can seamlessly integrate with central cloud resources, providing a balanced computational hierarchy. Furthermore, deploying AI at the network edge introduces new challenges in distributed learning and data privacy. The FL emerges as a viable solution, allowing AI models to be trained locally at the edge, with only model updates being shared to the central server. This reduces the bandwidth requirements and addresses privacy concerns by keeping sensitive data localized.
@@
The sustainability of 6G networks is a critical concern that must be addressed from an environmental and an economic perspective. AI has the potential to drive significant advancements in this area by optimizing network operations to reduce energy consumption.
Future research should focus on developing AI models that can accurately predict the energy consumption of network components and optimize their operation to minimize waste. This includes the dynamic adjustment of network infrastructure, such as BSs and data centers, based on user demand and traffic patterns. AI can enable these components to enter low-power states when demand is low and quickly ramp up when the demand increases. In addition to optimizing existing infrastructure, research should explore the design of new, energy-efficient hardware that can support AI operations at the network edge. This includes the development of specialized processors that can run AI algorithms with high efficiency and low power consumption.
Renewable energy sources in 6G networks can be another critical area of research. AI can manage integrating renewable energy sources, such as solar and wind, into the network’s power supply. This involves predicting energy generation patterns and adjusting network operations to match the availability of renewable energy. Furthermore, AI can contribute to developing new communication protocols designed with energy efficiency in mind. By optimizing how data is transmitted and received, these protocols can reduce the energy consumption of network devices and extend their operational lifetime.
@@
Ultra-low latency communications are a key feature of 6G networks, enabling a new generation of applications that require immediate response times. AI is poised to play a pivotal role in achieving this goal by providing real-time data processing and decision-making capabilities.
Future research should focus on developing AI systems that can operate at the network edge, providing near-instantaneous response times for critical applications. This includes creating lightweight AI models that can be deployed on edge devices with limited computational resources. One of the challenges in this area is the need for AI algorithms that can make decisions based on incomplete or uncertain data. Research should explore the use of techniques such as probabilistic modeling and approximate computing to enable AI systems to operate effectively under these conditions. Another critical research area can be the development of AI-driven network orchestration tools that can manage the allocation of computational resources in real time. These tools must be able to predict the computational needs of different applications and allocate resources accordingly to ensure that latency requirements are met.
@@
The security of 6G networks is of utmost importance, as these networks will support a wide range of critical applications and services. AI has the potential to significantly enhance the cybersecurity of 6G networks by providing advanced threat detection and response capabilities. Future research can look into developing AI-based security frameworks that can identify and respond to cyber threats in real time. This involves using machine learning algorithms to analyze network traffic and detect anomalies that may indicate a security breach.
One of the challenges in this area is the need for AI systems that can adapt to the constantly evolving landscape of cyber threats. Research should explore using adaptive learning algorithms that can update their models in response to new types of attacks. Another critical research area can be the development of AI-driven security protocols that can automatically patch vulnerabilities and respond to incidents without human intervention. This requires the creation of intelligent systems that can understand the context of security incidents and take appropriate actions to mitigate their impact. The use of AI for cybersecurity also raises significant privacy concerns. Research can address these concerns by developing privacy-preserving AI algorithms that protect user data while providing adequate security.
@@
The AI with 6G communication networks promises significant advancements in key technologies such as reconfigurable intelligence surfaces (RIS) [315], integrated sensing and communication (ISAC), Open Radio Access Networks (ORAN), Non-Terrestrial Networks (NTN) and near-field communication [316]. Intelligent reflecting surfaces can dynamically manipulate wireless environments to optimize signal propagation. Future research should develop advanced AI algorithms for real-time RIS optimization, leveraging machine learning to adapt to changing channel conditions. Additionally, the interaction between RIS and other 6G technologies, such as massive MIMO and millimeter-wave communications, needs thorough exploration to enhance spectral efficiency and network capacity.
The ISAC merges sensing and communication functions to improve spectrum utilization and provide environmental context for network optimization. AI-driven techniques for joint waveform design and resource allocation are crucial for seamless ISAC integration [317]. Moreover, developing AI algorithms for real-time data fusion will be essential for extracting actionable insights from ISAC systems. ORAN’s software-defined architecture can benefit from AI-driven network management, enhancing adaptability and resilience. NTN, including satellite systems, can leverage AI for efficient resource management and fault detection, ensuring robust global connectivity. Future research should focus on AI models tailored to NTNs, addressing challenges like latency and dynamic network topology [318]. By advancing AI applications in these key areas, researchers can unlock new levels of performance and efficiency in 6G networks, paving the way for a more connected and intelligent world.
@@ 10 Conclusions
In this all-encompassing review, we initially highlight that the 6G network’s evolution towards an integrated wireless infrastructure platform amalgamating communication, sensing, computation, intelligence, and storage hinges on the profound integration of AI and communications. The retrospective analysis of the development of AI and the evolution of large-scale models within the paper distinctly discloses its fundamental role in molding modern communication technology. Concurrently, the exposition of the three-stage integration of AI within 6G networks, namely “AI for Network,” “Network for AI,” and “AI as a Service,” furnishes a lucid blueprint for understanding the technological development trajectory. These three consecutive stages present the strategic shift from utilizing AI to optimize the network, through the network facilitating support for AI to the network offering AI as a service, thereby underlining the interdependent and coordinative developmental relationship between communication and AI. In studying the standardization process of AI within the network environment, we have systematically organized the key milestones and current priorities. This is vital for guaranteeing the coordinated global development of the technology. Analyzing the pros and cons of integrating AI and communication in 6G networks furnishes a foundation for a comprehensive evaluation of the technology’s impact. This enables us to maximize the benefits and minimize the risks during its development. Additionally, we have analyzed the challenges of integrating AI and communication from the perspectives of sustainability, real-time performance, security, etc. The proposed countermeasures for the current research challenges offer feasible directions for subsequent research and practice. We are convinced that this review holds profound significance for the research on the integration of AI and communication for the 6G network, and it can provide deep-seated insights for academic researchers and industry experts alike.

### v3

@@ 7.4 Requirement of Real-Time AI
The 6G aims to achieve uRLLC to support various real-time applications, including autonomous driving, telemedicine, and industrial automation. These scenarios impose unprecedented requirements on network real-time responsiveness and reliability, necessitating data transmission and processing to be completed within milliseconds or even microseconds. However, the current state-of-the-art AI models, characterized by large parameter spaces and intensive computation, typically employ distributed deployment, which inevitably introduces additional latency [283]. Furthermore, in data fusion scenarios that demand real-time data processing, traditional algorithms may encounter challenges in ensuring real-time performance. For example, in scenarios such as sensor data fusion in industrial automation and real-time traffic condition monitoring and analysis in intelligent transportation systems, algorithms are required to process data and make accurate fusion decisions within an extremely short time [284].
DL algorithms construct DNNs to automatically learn deep feature representations from vast and complex data, unlike traditional methods that rely on manually designed feature extraction rules. This significantly reduces errors and limitations caused by human factors, thereby enhancing the accuracy and real-time performance of data fusion. RL dynamically adjusts fusion weights and methods based on the real-time feedback information from different data sources to adapt to the constantly changing data environment and task requirements. Reducing the computational requirements of models without significantly compromising accuracy is possible through techniques such as model pruning, parameter quantization, and adopting more efficient network architectures, such as lightweight versions of CNN and RNN [285]. Additionally, adopting edge computing can reduce latency by shifting data processing tasks from the central cloud to the network edge [286]. Innovations in network architecture, such as network slicing technology, can also provide customized network resources for different services and applications, ensuring critical tasks receive the necessary bandwidth and latency guarantees [287]. Moreover, AI can manage network resources dynamically and accommodate real-time changes in network conditions. The development of hardware acceleration is also crucial for achieving real-time AI. Dedicated AI processors, such as GPU, Tensor Processing Units, and Field-Programmable Gate Arrays, can expedite the inference process of models, helping to meet the stringent real-time requirements of 6G networks.
@@ 7.5 Security and Privacy
The integration of AI and 6G introduces complex security and privacy challenges. AI algorithms are susceptible to adversarial attacks, where slight, often imperceptible, perturbations in the input data can lead to incorrect outputs, thereby compromising the system’s integrity. Moreover, the security and privacy challenges faced by AI algorithms in the process of data fusion are extremely formidable. Since the data being processed generally originates from numerous diverse data sources, which frequently encompass sensitive information such as personal identities, medical records, and financial transactions. This thereby gives rise to risks prevailing in all aspects of data fusion, ranging from the storage and transmission of data to its processing [288]. For instance, the server where the fused data is stored might be subject to hacker attacks, subsequently leading to data leakage. Moreover, during the data preprocessing stage of the AI models training, if appropriate encryption protection measures are lacking, the data is highly susceptible to being stolen.
Some existing work has inspired addressing these security and privacy issues. One aspect involves the development of robust AI models through techniques like adversarial training, which can enhance the resilience of AI systems against malicious to mitigate these inputs [289]. Another approach is the implementation of FL, where AI models are trained across multiple decentralized devices without exchanging the raw data, thereby preserving privacy [290]. Moreover, the authors in [291] proposed an adaptive feature-correlation region segmentation mechanism, which maintained a sound privacy protection for the DL model under a given moderate privacy budget. In addition to technological solutions, regulatory frameworks and policies must be established to define clear data protection and cybersecurity standards in the context of AI-6G integration. This includes setting guidelines for responsible AI usage, ensuring compliance with privacy regulations such as the general data protection regulation, and fostering international collaboration to combat cyber threats. Moreover, SMPC and homomorphic encryption can perform computations on encrypted data, ensuring data privacy even when processed by third parties [292].
@@ 7.6 Human-AI Interactive Collaboration
Integrating AI with 6G communication networks presents transformative opportunities for enhancing human-AI interactive collaboration. However, this convergence also introduces significant challenges that must be addressed to realize its full potential. One of the primary issues is the stringent real-time response requirements imposed by 6G’s ultra-low latency capabilities. conventional AI algorithms, which often require substantial computational resources and time, may struggle to develop this, so developing lightweight AI models optimized for edge computing is essential. These models must be capable of rapid processing and decision-making, reducing the dependency on centralized data centers and minimizing latency. Furthermore, 6G networks will facilitate the transmission of diverse data types, including video, audio, text, and sensor data, necessitating human-AI interactive collaboration systems to handle and interpret multimodal data adeptly. Advances in DL and multimodal learning technologies are crucial for effectively fusing and understanding these varied data streams, thereby enhancing the naturalness and intuitiveness of interactions.
@@ 8 Future Works
The advent of 6G networks heralds a paradigm shift in managing network operations, with AI playing a central role. Adaptive learning mechanisms, leveraging AI, are set to become the cornerstone of smart network management. These mechanisms must handle the complexity and dynamism of 6G environments characterized by heterogeneous devices, variable traffic patterns, and the need for real-time processing. Research is expected to focus on developing ML models that can continuously learn and adapt without human intervention. These models should be scalable to manage the exponential increase in connected devices and robust to ensure consistent performance under varying network states. One promising approach is using DRL, which can dynamically optimize network decisions. DRL can adjust network parameters in real-time, ensuring optimal performance for changing network loads and user mobility patterns. Moreover, integrating transfer learning techniques can enable these AI models to apply knowledge gained in one domain to new but related problems. This is particularly relevant in 6G networks that will operate in diverse environments and must quickly adapt to new conditions. Transfer learning can reduce the need for extensive retraining of models, thereby saving computational resources and time. Another focus area can be the development of self-healing networks using AI. By incorporating predictive analytics, networks can anticipate failures and automatically reroute traffic or adjust parameters to mitigate the impact. This proactive approach to fault management will be critical for maintaining the high reliability and availability required by 6G applications.
@@
The underlying network must be optimized to cater to their requirements for AI applications to reach their full potential. 6G networks, with their promise of higher bandwidths and lower latencies, provide an ideal platform for AI applications but also pose significant challenges in terms of architectural design. Future research can delve into creating network architectures that are inherently flexible and can dynamically adapt to the needs of AI applications. Such architectures must support high data rates and the ability to process large volumes of data with minimal latency. AI can significantly enhance the capability of the network to manage these requirements by predicting traffic patterns and adjusting resources accordingly. A key component of such architectures is edge computing, which brings computational resources closer to the data sources. This is crucial for latency-sensitive AI applications that require near-instantaneous decision-making. Research should explore the development of edge computing solutions that can seamlessly integrate with central cloud resources, providing a balanced computational hierarchy. Furthermore, deploying AI at the network edge introduces new challenges in distributed learning and data privacy. The FL emerges as a viable solution, allowing AI models to be trained locally at the edge, with only model updates being shared to the central server. This reduces the bandwidth requirements and addresses privacy concerns by keeping sensitive data localized.
@@
The sustainability of 6G networks is a critical concern that must be addressed from an environmental and an economic perspective. AI has the potential to drive significant advancements in this area by optimizing network operations to reduce energy consumption. Future research should focus on developing AI models that can accurately predict the energy consumption of network components and optimize their operation to minimize waste. This includes the dynamic adjustment of network infrastructure, such as BSs and data centers, based on user demand and traffic patterns. AI can enable these components to enter low-power states when demand is low and quickly ramp up when the demand increases. In addition to optimizing existing infrastructure, research should explore the design of new, energy-efficient hardware that can support AI operations at the network edge. This includes the development of specialized processors that can run AI algorithms with high efficiency and low power consumption. Renewable energy sources in 6G networks can be another critical area of research. AI can manage integrating renewable energy sources, such as solar and wind, into the network’s power supply. This involves predicting energy generation patterns and adjusting network operations to match the availability of renewable energy. Furthermore, AI can contribute to developing new communication protocols designed with energy efficiency in mind. By optimizing how data is transmitted and received, these protocols can reduce the energy consumption of network devices and extend their operational lifetime.
@@
Ultra-low latency communications are a key feature of 6G networks, enabling a new generation of applications that require immediate response times. AI is poised to play a pivotal role in achieving this goal by providing real-time data processing and decision-making capabilities. Future research should focus on developing AI systems that can operate at the network edge, providing near-instantaneous response times for critical applications. This includes creating lightweight AI models that can be deployed on edge devices with limited computational resources. One of the challenges in this area is the need for AI algorithms that can make decisions based on incomplete or uncertain data. Research should explore the use of techniques such as probabilistic modeling and approximate computing to enable AI systems to operate effectively under these conditions. Another critical research area can be the development of AI-driven network orchestration tools that can manage the allocation of computational resources in real time. These tools must be able to predict the computational needs of different applications and allocate resources accordingly to ensure that latency requirements are met.
@@
The security of 6G networks is of utmost importance, as these networks will support a wide range of critical applications and services. AI has the potential to significantly enhance the cybersecurity of 6G networks by providing advanced threat detection and response capabilities. Future research can look into developing AI-based security frameworks that can identify and respond to cyber threats in real time. This involves using ML algorithms to analyze network traffic and detect anomalies that may indicate a security breach. One of the challenges in this area is the need for AI systems that can adapt to the constantly evolving landscape of cyber threats. Research should explore using adaptive learning algorithms that can update their models in response to new types of attacks. Another critical research area can be the development of AI-driven security protocols that can automatically patch vulnerabilities and respond to incidents without human intervention. This requires the creation of intelligent systems that can understand the context of security incidents and take appropriate actions to mitigate their impact. The use of AI for cybersecurity also raises significant privacy concerns. Research can address these concerns by developing privacy-preserving AI algorithms that protect user data while providing adequate security.
@@
The AI with 6G communication networks promises significant advancements in key technologies such as intelligent reflecting surface (IRS) [293], integrated sensing and communication (ISAC), O-RAN, non-terrestrial networks (NTN) and near-field communication [294]. For example, IRS can dynamically manipulate wireless environments to optimize signal propagation. Future research should develop advanced AI algorithms for high dimensional and real-time IRS beamforming and channel estimation, leveraging ML to adapt to changing channel conditions in practical complex scenarios. The ISAC merges sensing and communication functions to improve spectrum utilization and provide environmental context for network optimization. AI-driven techniques for joint waveform design and resource allocation are crucial for achieving seamless ISAC integration [295]. Moreover, developing AI algorithms for real-time data fusion will be essential for extracting actionable insights from ISAC systems. ORAN’s software-defined architecture can benefit from AI-driven network management, enhancing adaptability and resilience. NTN, including satellite systems, can leverage AI for efficient resource management and fault detection, ensuring robust global connectivity. Future research should focus on AI models tailored to NTNs, uniquely addressing challenges like latency and dynamic network topology. By advancing AI applications in these key areas, researchers can unlock new levels of performance and efficiency in 6G networks, paving the way for a more connected and intelligent world.
Ubiquitous Computing
With the proliferation of intelligent devices and the advancement of IoT technologies, the demand for intelligent services from users and devices across wide areas continues to grow. Consequently, the integration of 6G and AI aims to provide seamless AI services to users and devices worldwide, achieving ubiquitous computing and meeting the diverse intelligent service needs of users and devices. By deploying edge servers on unmanned aerial vehicle (UAV) and low earth orbit (LEO) satellites to implement MEC, multi-layered and heterogeneous computing services can be delivered to users and devices, satisfying their differentiated QoS requirements. The primary challenge in this domain lies in designing rational, efficient, and dynamic resource allocation schemes for heterogeneous networks that integrate terrestrial BSs, UAVs, and LEO satellites. To address this, research should be based on the characteristics of integrated terrestrial, aerial, and satellite network architectures, leveraging theories such as convex optimization and Lyapunov optimization. Particular emphasis should be placed on advanced algorithms like DRL to explore computation offloading strategies and resource allocation scheme designs.
@@ 9 Conclusions
In this all-encompassing review, we have initially highlighted that the 6G network’s evolution towards an integrated wireless infrastructure platform amalgamating communication, sensing, computation, intelligence, and storage hinges on the profound integration of AI and communications. Concurrently, the exposition of the three-stage integration of AI within 6G networks, namely “AI4NET,” “NET4AI,” and “AIaaS,” furnishes a lucid blueprint for understanding the technological development trajectory. These three consecutive stages present the strategic shift from utilizing AI to optimize the network, through the network facilitating support for AI to the network offering AIaaS, thereby underlining the interdependent and coordinative developmental relationship between communication and AI. The 6G networks will introduce entirely new service paradigms, seamlessly integrating AI capabilities into various application scenarios, thereby serving as a powerful platform to support the advancement of AI In studying the standardization process of AI for wireless networks, we have emphasized the crucial milestones and the ongoing efforts. We have further explored the challenges faced by the integration of AI and communications in the 6G era, including considerations of sustainability, the demand for real-time AI, issues of security and privacy, as well as human-AI interaction and collaboration. Finally, we have outlined promising future research opportunities that are expected to advance the development and optimization of AI and 6G communications. We are convinced that this review holds profound significance for the research on the integration of AI and communication for the 6G network, and it can provide deep-seated insights for academic researchers and industry experts alike.
@@
Appendix A
Table 4: A list of abbreviations
Abbreviation Definition Abbreviation Definition
3GPP 3rd generation partnership project LLM large language model
4G fourth generation LoS line of sight
5G fifth generation LAM large-scale AI model
6G sixth generation ML machine learning
AI artificial intelligence MIMO multiple-input multiple-output
AI4NET AI for network mMTC massive machine type communications
AIaaS AI as a service MEC mobile edge computing
AR augmented reality MARL multi-agent reinforcement learning
BS base station NLP natural language processing
CSI channel state information NLoS non-line of sight
CNN convolutional neural networks NWDAF network data analytics function
CN core network NTN non-terrestrial network
CP cyclic prefix NR new radio
CPU central processing units NET4AI network for AI
DL deep learning O&M operation and maintenance
DRL deep reinforcement learning OFDM orthogonal frequency division multiplexing
DT digital twins O-RAN open radio access networks
DQN deep Q-network PFL personalized federation learning
DDPG deep deterministic policy gradient QoS quality of service
DNN deep neural network QoAIS quality of AI service
eMBB enhanced mobile broadband RL reinforcement learning
ETSI european telecommunication standardization institute RSS received signal strength
ENI experiential networked intelligence RAN radio access network
FL federated learning REK radio environment knowledge
FedAvg federated averaging RNN recurrent neural network
gNB generation node B SGCS square of the generalized cosine similarity
GPS global positioning system SNR signal-to-noise ratio
GPU graphics processing units SL split learning
(Continued on next page)
Abbreviation Definition Abbreviation Definition
HARQ hybrid automatic repeat request SMPC secure multi-party computation
ITU international telecommunication union SI semantic information
ICT information and communication technology SA system aspects
INC In-network computing TSG technical specification groups
IP internet protocol uRLLC ultra-reliable low latency communications
IMT international mobile communications UPF user plane functions
IRS intelligent reflecting surface UAV unmanned aerial vehicle
ISAC integrated sensing and communication VR virtual reality
IoT internet of things WG working groups
KPI key performance indicator XR extensive reality
LEO low earth oribit